OpenStudy (anonymous):

1/sqrt(x) = sqrt(2) / 2

OpenStudy (anonymous):

yes it is true, since \[\frac{1}{\sqrt{x}}=\frac{1}{\sqrt{x}}\times \frac{\sqrt{x}}{\sqrt{x}}=\frac{\sqrt{x}}{x}\]